# How Numerical Precision Affects Mathematical Reasoning Capabilities of LLMs

### v1

@@ Appendix E Proofs for Section 5
For any integers $n$ and $p$, there exists a logarithmic-precision Transformer with constant depth $L$ and constant hidden dimension $d$ (independent of $n$) that can generate the correct output for any input on the $\operatorname{ADD}(n,p)$ task.
@@
Theorem 5.1 is a special case of Theorem 5.2. Taking $k=2$ in Theorem 5.2 completes our proof. Notice that $m=\lceil\log_{p}k\rceil=1$, which implies that we don’t need to combine neighboring bits actually. ∎
@@
For any integers $n$, $k$, and $p$, there exists a logarithmic-precision Transformer with constant depth $L$ and constant hidden dimension $d$ (independent of $n$ and $k$) that can generate the correct output for any input on the $\operatorname{IterADD}(n,k,p)$ task.
For ease of reading, we also describe an algorithm to perform $\operatorname{IterADD}(n,k,p)$ (Algorithm 2) and prove the correctness first. Then, we will construct a constant-size Transformer with logarithmic precision to simulate Algorithm 2.
Input :  $k$ $p$-adic numbers ${\bm{a}}_{1},\cdots,{\bm{a}}_{k}$, with maximum length $n$
@@ 1 $m=\lceil\log_{p}k\rceil$;
2 Compute the sum of each bit as $r_{j}=\sum_{i\in[k]}a_{ij}$ for $j=0,\cdots,n-1$;
@@ 5 $c_{0}=0$;
7       Compute the carry-on bits ${\bm{c}}$:
@@ 14 Covert $p^{m}$-adic $\tilde{\bm{o}}$ to $p$-adic ${\bm{o}}$:
for $i=jk+l$ where $l\in\{0,\cdots,k-1\}$, $j\in{\mathbb{Z}}$;
Algorithm 2 $\operatorname{IterADD}(n,k,p)$ Algorithm
Lemma E.1 (An algorithm to perform $\operatorname{IterADD}(n,k,p)$).
Algorithm 2 outputs ${\bm{o}}={\bm{a}}_{1}+\cdots+{\bm{a}}_{k}$ for all inputs ${\bm{a}}_{1},\cdots,{\bm{a}}_{k}$.
@@
In the initial four steps, we convert $p$-adic addition into $p^{m}$-adic addition. This transformation enables the expression $\sum_{i}s_{i}p^{im}$ to accurately represent the sum of $k$ numbers.
At this stage, $s_{i}\in[0,kp^{m})$. To compute the final results accurately, it is crucial to account for carry-over effects to ensure the final outputs $\tilde{o}_{i}$ range from 00 to $p^{m}-1$. We can decompose each $s_{i}$ using the formula $s_{i}=b_{i}p^{m}+q_{i}$, where $q_{i}\in[0,p^{m}-1]$ and $b_{i}<k\leq p^{m}$. Consequently, the overflow $b_{i}$ will only directly overflow to the immediately subsequent bit $q_{i+1}$. And we have $q_{i+1}+b_{i}\leq 2(p^{m}-1)$. Therefore, let ${\bm{c}}$ denote the vector to record the carry-over effects on the position $i$, and we have
@@
$c_{i}=(c_{i-1}\cdot{\bm{1}}_{q_{i}+b_{i-1}\geq p^{m}-1})\vee{\bm{1}}_{q_{i}+b_{i-1}\geq p^{m}}.$
By removing the recursive computation, we can compute the carry-over by the following formulas:
@@
In Equation 8, $i_{\wedge}$ indicates the nearest bit that can contribute an additional carry to the higher bits up to the $i$-th bit, while $i_{\vee}$ identifies the highest bit lower than $i$, such that the carry generated by bits below $i_{\vee}$ does not affect bits higher than $i_{\vee}$. Therefore, the carry-over $c_{i}=1$ if and only if $i_{\wedge}>i_{\vee}$.
Then after computing the carry-over vectors, we can compute the $p^{m}$-adic sum of the input integers as follows:
@@
Finally, to convert the $p^{m}$-adic number to the final $p$-adic number $\tilde{{\bm{o}}}$, we only need to perform the modulus operation:
for $i=jk+l$ where $l\in\{0,\cdots,k-1\}$, $j\in{\mathbb{Z}}$. Therefore, the output ${{\bm{o}}}$ is exactly the sum of the $m$ $n$-bits integers and the Algorithm 2 outputs $\operatorname{IterADD}({\bm{a}}_{1},\cdots,{\bm{a}}_{m})$ for all ${\bm{a}}_{1},\cdots,{\bm{a}}_{m}$. ∎
Next, we provide the proof for Theorem 5.2.
Proof for Theorem 5.2.
Now, we demonstrate that a log-precision transformer, with a constant depth, a fixed number of attention heads, and constant embedding dimensions, is capable of simulating Algorithm 2. Consequently, this model can accurately generate correct output for any input integers ${\bm{a}}_{1},\dots,{\bm{a}}_{k}$.
Initial Embeddings: The total length of the input sequence is no longer than $k(n+1)$. We categorize the tokens into two classes: number tokens ($10,1,\cdots,p-$ 1) and auxiliary tokens ($+$, $=$, <SOS> and <EOS>). Given the parameters $k,n$, we determine the parameter $m=\lceil\log_{p}k\rceil$, as specified in Algorithm 2. The embeddings for these classes are defined as follows:
  • Embedding of input token $a_{i,j}$: ${\bm{e}}_{i,j}^{0}=\left(a_{i,j},0,0,i,j,j\operatorname{mod}m,\lfloor\frac{j}{m}\rfloor,p^{j\operatorname{mod}m},p^{-(j\operatorname{mod}m)},\text{ape}_{i,j}\right)$.
@@
  • Embedding of the <SOS> token: ${\bm{e}}_{\texttt{<SOS>}}^{0}=(0,1,0,0,-1,-1,-1,0,0,\text{ape}_{\texttt{<SOS>}})$.
  • Embedding of the <EOS> token: ${\bm{e}}_{\texttt{<EOS>}}^{0}=(0,0,1,0,-1,-1,-1,0,0,\text{ape}_{\texttt{<EOS>}})$.
  • Embedding of output token $o_{i}$: ${\bm{e}}_{o,i}^{0}=({o}_{i},0,0,0,i,i\operatorname{mod}m,\lfloor\frac{i}{m}\rfloor,p^{i\operatorname{mod}m},p^{-(i\operatorname{mod}m)},\text{ape}_{o,i})$.
where $\text{ape}_{\cdots}$ is the absolute positional encoding. In this construction, the first three dimensions of each initial embedding represent the word embedding, while the last six dimensions account for the position embedding.
Block 1. The first block of the Transformer calculates the following values:
@@
    $l_{i,j}$: the number of previous tokens (inclusive) $a_{i^{\prime},j^{\prime}}$ where $i^{\prime}=i$ and $\lfloor\frac{j^{\prime}}{m}\rfloor=\lfloor\frac{j}{m}\rfloor$. $l_{i,j}$ is only defined on input number tokens $a_{i,j}$. If the value is undefined, we will set $l=-1$.
@@
    $f_{i,j}$: $f_{i,j}=1$ if no previous tokens (exclusive) $a_{i^{\prime},j^{\prime}}$ with $\lfloor\frac{j^{\prime}}{m}\rfloor=\lfloor\frac{j}{m}\rfloor$. $f_{i,j}$ is only defined on input number tokens $a_{i,j}$. If the value is undefined, we will set $f=-1$.
To calculate the first value, we can set the query, key, value and $r$ in Section C.2 as follows:
@@
We can calculate the values required in query or key by previous MLPs using Lemma C.1. Thus $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=-\left(\lfloor\frac{j^{\prime}}{m}\rfloor-\lfloor\frac{j}{m}\rfloor\right)^{2}-(i^{\prime}-i)^{2}$, which implies that $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=0$ if $\lfloor\frac{j^{\prime}}{m}\rfloor=\lfloor\frac{j}{m}\rfloor$ and $i=i^{\prime}$, and $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle\leq-1$ otherwise. By Lemma C.5, we can use one attention head to copy the absolute position $j^{\prime\prime}$ of the first token satisfying the conditions. Finally, the number of the tokens is $j-j^{\prime\prime}+1$.
To calculate the second value, we can set the query, key, and $r$ in Section C.2 as follows:
@@
Thus $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=-\left(\lfloor\frac{j^{\prime}}{m}\rfloor-\lfloor\frac{j}{m}\rfloor\right)^{2}$, which implies that $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=0$ if $\lfloor\frac{j^{\prime}}{m}\rfloor=\lfloor\frac{j}{m}\rfloor$, and $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle\leq-1$ otherwise. By Lemma C.5, we can use one attention head to copy the absolute position $j^{\prime\prime}$ of the first token satisfying the conditions. Finally, we have $f_{i,j}=1$ if $j^{\prime\prime}=j$. It suffices to determine whether $j^{\prime\prime}=j$. Notice that ${\bm{1}}_{j^{\prime\prime}=j}=\text{ReLU}[1-(j-j^{\prime\prime})^{2}]$, thus we can calculate the value of $f_{i,j}$ by a constant-width MLP using Lemma C.1.
Finally, if the values are undefined, those values will be set as $-1$ in the MLP stage by using conditional selection (Lemma C.3) using information in positional embeddings. To sum up, the new embeddings generated in this block can be summarized as: ${\bm{e}}^{1}=(l,f)$. All these embeddings will be concatenated with the original embeddings.
Block 2. The second block of the Transformer is tailored to execute the first three lines of Algorithm 2. In this block, the model aggregates the adjacent $m$ bits to derive $s_{i}$ via the attention mechanism. For each token $a_{i,j}$, denote $t_{i,j}$ as the number of previous tokens (inclusive) $a_{i^{\prime},j^{\prime}}$ where $\lfloor\frac{j}{m}\rfloor=\lfloor\frac{j^{\prime}}{m}\rfloor$. The second block of the Transformer calculates the following values:
@@
    $\frac{1}{t_{i,j}}$ where $t_{i,j}$ is defined above. If the value is undefined, we will set the value as $-1$.
@@
    $c_{i,j}$: the mean value applied across $a_{i^{\prime},j^{\prime}}p^{j^{\prime}\operatorname{mod}m}$ for previous tokens (inclusive) $a_{i^{\prime},j^{\prime}}$ where $\lfloor\frac{j^{\prime}}{m}\rfloor=\lfloor\frac{j}{m}\rfloor$. If the value is undefined, we will set the value as $-1$.
Then we can get $s_{w}=c_{i,mw}t_{i,mw}$ where $i$ is the maximize value such that the length of ${\bm{a}}_{i}$ is greater than $mk$.
To calculate the first value, we first set the query, key and value as follows:
@@
These values can be calculated using Lemma C.1. Thus $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=-\left(\lfloor\frac{j^{\prime}}{m}\rfloor-\lfloor\frac{j}{m}\rfloor\right)^{2}$, which implies that $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=0$ if $\lfloor\frac{j^{\prime}}{m}\rfloor=\lfloor\frac{j}{m}\rfloor$, and $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle\leq-1$ otherwise. By the definition of $f_{i,j}$ and Lemma C.6, we can get the attention output is $\frac{1}{t_{i,j}}$, as desired.
To calculate the second value, we can set the query, key and value as follows:
@@
Similar to the calculation of the first value, these values can be calculated using Lemma C.1. Thus $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=-\left(\lfloor\frac{j^{\prime}}{m}\rfloor-\lfloor\frac{j}{m}\rfloor\right)^{2}$, which implies that $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=0$ if $\lfloor\frac{j^{\prime}}{m}\rfloor=\lfloor\frac{j}{m}\rfloor$, and $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle\leq-1$ otherwise. By Lemma C.6, the attention output is $c_{i,j}$, as desired.
Finally, if the values are undefined, those values will be set as $-1$ in the MLP stage by using conditional selection (Lemma C.3) using information in positional embeddings. To sum up, the new embeddings generated in this block can be summarized as: ${\bm{e}}^{2}=(\frac{1}{t},c)$. All these embeddings will be concatenated with the original embeddings.
@@
Block 3. The third block of the Transformer calculates the value of $c_{i,j}t_{i,j}$. We will first calculate $t_{i,j}$ by attention layer and $\frac{1}{t_{i,j}}$ from the last block, then calculate $c_{i,j}t_{i,j}$ using Lemma C.1.
@@
Notice that $t_{i,j}$ will not exceed the absolute positional of the current token. We set the query, key and value as follows:
  • Query: ${\bm{e}}_{i,j}=\left(\frac{1}{t_{i,j}^{2}},-\frac{2}{t_{i,j}},1\right)$.
  • Key: ${\bm{k}}_{i^{\prime},j^{\prime}}=\left(\text{ape}_{i^{\prime},j^{\prime}}^{2},\text{ape}_{i^{\prime},j^{\prime}},1\right)$.
  • Value: ${\bm{v}}_{i^{\prime},j^{\prime}}=(\text{ape}_{i^{\prime},j^{\prime}})$.
@@
These values can be calculated using Lemma C.1. Thus $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=-\left(\frac{\text{ape}_{i^{\prime},j^{\prime}}}{t_{i,j}}-1\right)^{2}$, which implies that $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=0$ if $\text{ape}_{i^{\prime},j^{\prime}}=t_{i,j}$, and $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle\leq-\frac{1}{n^{2}k^{2}}$ otherwise since $t_{i,j}\leq nk$. By Lemma C.6, the attention output is $t_{i,j}$, as desired.
@@
Finally, we can calculate the value of $c_{i,j}t_{i,j}$ using the subsequent MLP. To sum up, the new embeddings generated in this block can be summarized as: ${\bm{e}}^{3}=(ct)$. All these embeddings will be concatenated with the original embeddings.

### v2

@@ Appendix E Proofs for Section 5
In this section, we provide the formal proofs of the theorems stated in Section 5. Before proceeding with the proofs, we revisit the role of the tokenizer ${\bm{T}}_{c}$. As established in Section B.4 and Proposition B.6, it is sufficient to focus on the case of ${\bm{T}}_{1}$, where both the input and output sequences are tokenized into single digits. This simplification is crucial for the subsequent analysis and constructions. Notably, this reasoning parallels the argument presented at the beginning of Appendix D (Proof of Section 4).
E.1 Proof for Theorem 5.1
@@
Fix integers $p\geq 2$ and $c\in\mathbb{N}^{*}$. Consider the tokenizer ${\bm{T}}_{c}$ defined in Eq. 1 for processing the input and output sequences. There exists a logarithmic-precision Transformer with constant depth and hidden dimension (independent of $n$) that can generate the correct output for any input on the $\operatorname{ADD}_{p}(n)$ task.
@@
The result in Theorem 5.1 follows as a special case of Theorem 5.2. Specifically, by setting $k=2$ in Theorem 5.2, the proof is complete. Observe that in this case, $m=\lceil\log_{p}k\rceil=1$, which implies that the combination of neighboring bits is unnecessary. ∎
E.2 Proof for Theorem 5.2
@@
Fix integers $p\geq 2$ and $c\in\mathbb{N}^{*}$. Consider the tokenizer ${\bm{T}}_{c}$ defined in Eq. 1 for processing the input and output sequences. For any integers $n$ and $k$, there exists a logarithmic-precision Transformer with constant depth and hidden dimension $d$ (independent of $n$ and $k$) that can generate the correct output for any input on the $\operatorname{IterADD}_{p}(n,k)$ task.
For ease of understanding, we first present an algorithm to compute $\operatorname{IterADD}_{p}(n,k)$ (Algorithm 2) and prove its correctness. Subsequently, we demonstrate the construction of a constant-size Transformer with logarithmic precision to simulate Algorithm 2.
Input :  $k$ $p$-adic numbers ${\bm{a}}_{1},\cdots,{\bm{a}}_{k}$, each of maximum length $n$
@@ 1 $m=\lceil\log_{p}k\rceil$;
2 Compute the sum of each bit: $r_{j}=\sum_{i\in[k]}a_{ij}$ for $j=0,\cdots,n-1$;
@@ 5 Initialize $c_{0}=0$;
7       Compute carry bits ${\bm{c}}$:
@@ 14 Covert $p^{m}$-adic $\tilde{\bm{o}}$ to $p$-adic ${\bm{o}}$:
for $i=jk+l$, where $l\in\{0,\cdots,k-1\}$, $j\in{\mathbb{Z}}$;
Algorithm 2 $\operatorname{IterADD}_{p}(n,k)$ Algorithm
Lemma E.1 (Algorithm for $\operatorname{IterADD}_{p}(n,k)$).
Algorithm 2 computes ${\bm{o}}={\bm{a}}_{1}+\cdots+{\bm{a}}_{k}$ for any inputs ${\bm{a}}_{1},\ldots,{\bm{a}}_{k}$.
@@
The initial four steps of the algorithm transform $p$-adic addition into $p^{m}$-adic addition. This transformation allows the sum of $k$ numbers to be represented as $\sum_{i}s_{i}p^{im}$, where $s_{i}$ are intermediate coefficients.
At this stage, $s_{i}\in[0,kp^{m})$. To ensure the final results are accurate, we must account for carry-over effects such that the outputs $\tilde{o}_{i}$ remain within the range $[0,p^{m}-1]$. Each $s_{i}$ can be decomposed as $s_{i}=b_{i}p^{m}+q_{i}$, where $q_{i}\in[0,p^{m}-1]$ and $b_{i}<k\leq p^{m}$. Consequently, the overflow $b_{i}$ propagates only directly to the next subsequent digit $q_{i+1}$. Notably, $q_{i+1}+b_{i}\leq 2(p^{m}-1)$.
Let ${\bm{c}}$ denote the vector recording carry-over effects at each position $i$. The carry-over can be computed iteratively as:
@@
$c_{i}=\big{(}c_{i-1}\cdot{\bm{1}}_{q_{i}+b_{i-1}\geq p^{m}-1}\big{)}\vee{\bm{1}}_{q_{i}+b_{i-1}\geq p^{m}}.$
To avoid recursive computation, the carry-over can also be derived using:
@@
Here, $i_{\wedge}$ identifies the highest bit contributing a carry to the $i$-th position, while $i_{\vee}$ identifies the highest bit below $i$ such that carry propagation from bits below $i_{\vee}$ does not affect higher bits. Thus, $c_{i}=1$ if and only if $i_{\wedge}>i_{\vee}$.
Once the carry-over vector ${\bm{c}}$ is determined, the $p^{m}$-adic sum can be computed as:
@@
Finally, to convert the $p^{m}$-adic representation back to a $p$-adic number $\tilde{{\bm{o}}}$, we perform the following modulus operation:
$o_{i}=\left\lfloor\frac{\tilde{o}_{j}\operatorname{mod}p^{(l+1)}}{p^{l}}\right\rfloor,$
for $i=jk+l$, where $l\in\{0,\dots,k-1\}$ and $j\in\mathbb{Z}$.
Therefore, the output ${\bm{o}}$ is precisely the sum of the $k$ input $p$-adic numbers, and the algorithm in Algorithm 2 correctly computes $\operatorname{IterADD}_{p}({\bm{a}}_{1},\ldots,{\bm{a}}_{k})$ for all ${\bm{a}}_{1},\ldots,{\bm{a}}_{k}$. ∎
Now we present the proof of Theorem 5.2.
Proof of Theorem 5.2.
We demonstrate that a log-precision transformer, with constant depth, a fixed number of attention heads, and constant embedding dimensions, is capable of simulating Algorithm 2. As a result, this model can reliably produce the correct output for any input integers ${\bm{a}}_{1},\dots,{\bm{a}}_{k}$.
Initial Embeddings: The total length of the input sequence is at most $k(n+1)$. Tokens in the sequence are divided into two categories: numeric tokens ($10,1,\dots,p-$ 1) and auxiliary tokens ($+$, $=$, <SOS>, <EOS>). Given the parameters $k$ and $n$, we define the parameter $m=\lceil\log_{p}k\rceil$, as specified in Algorithm 2. The embeddings for these tokens are constructed as follows:
  • Embedding of input token $a_{i,j}$:
    ${\bm{e}}_{i,j}^{0}=\left(a_{i,j},0,0,i,j,j\operatorname{mod}m,\lfloor\frac{j}{m}\rfloor,p^{j\operatorname{mod}m},p^{-(j\operatorname{mod}m)},\text{ape}_{i,j}\right).$
@@
    ${\bm{e}}_{=}^{0}=\left(0,1,0,k+1,-1,-1,-1,0,0,\text{ape}_{=}\right).$
  • Embedding of the <SOS> token:
    ${\bm{e}}_{\texttt{<SOS>}}^{0}=\left(0,1,0,0,-1,-1,-1,0,0,\text{ape}_{\texttt{<SOS>}}\right).$
  • Embedding of the <EOS> token:
    ${\bm{e}}_{\texttt{<EOS>}}^{0}=\left(0,0,1,0,-1,-1,-1,0,0,\text{ape}_{\texttt{<EOS>}}\right).$
  • Embedding of output token $o_{i}$:
    ${\bm{e}}_{o,i}^{0}=\left(o_{i},0,0,0,i,i\operatorname{mod}m,\lfloor\frac{i}{m}\rfloor,p^{i\operatorname{mod}m},p^{-(i\operatorname{mod}m)},\text{ape}_{o,i}\right).$
Here, $\text{ape}_{\cdots}$ represents the absolute positional encoding. In this construction, the first three dimensions of each embedding correspond to the word embedding, while the remaining six dimensions capture the positional embedding.
Block 1. The first block of the Transformer computes the following quantities:
@@
    $l_{i,j}$: The number of preceding tokens (inclusive) $a_{i^{\prime},j^{\prime}}$ satisfying $i^{\prime}=i$ and $\lfloor\frac{j^{\prime}}{m}\rfloor=\lfloor\frac{j}{m}\rfloor$. The value $l_{i,j}$ is defined only for input number tokens $a_{i,j}$. If undefined, we set $l=-1$.
@@
    $f_{i,j}$: Defined as $f_{i,j}=1$ if no preceding tokens (exclusive) $a_{i^{\prime},j^{\prime}}$ exist such that $\lfloor\frac{j^{\prime}}{m}\rfloor=\lfloor\frac{j}{m}\rfloor$; otherwise, $f_{i,j}=0$. This value is defined only for input number tokens $a_{i,j}$, and if undefined, we set $f=-1$.
To compute $l_{i,j}$, we define the query, key, value, and $r$ in Section C.2 as follows:
@@
Using Lemma C.1, the components of the query and key can be computed by preceding MLP layers. The result of the dot product is given by:
$\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=-\left(\lfloor\frac{j^{\prime}}{m}\rfloor-\lfloor\frac{j}{m}\rfloor\right)^{2}-(i^{\prime}-i)^{2}.$
This implies that $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=0$ if $\lfloor\frac{j^{\prime}}{m}\rfloor=\lfloor\frac{j}{m}\rfloor$ and $i=i^{\prime}$, and $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle\leq-1$ otherwise. By Lemma C.5, one attention head can be used to copy the absolute position $j^{\prime\prime}$ of the first token satisfying these conditions. The value of $l_{i,j}$ is then computed as $j-j^{\prime\prime}+1$.
To compute $f_{i,j}$, we redefine the query, key, and $r$ as follows:
@@
In this case, the dot product is given by:
$\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=-\left(\lfloor\frac{j^{\prime}}{m}\rfloor-\lfloor\frac{j}{m}\rfloor\right)^{2}.$
This yields $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=0$ if $\lfloor\frac{j^{\prime}}{m}\rfloor=\lfloor\frac{j}{m}\rfloor$, and $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle\leq-1$ otherwise. By Lemma C.5, one attention head can copy the absolute position $j^{\prime\prime}$ of the first token satisfying these conditions. The value $f_{i,j}$ is then determined by checking whether $j^{\prime\prime}=j$. Specifically, we evaluate:
${\bm{1}}_{j^{\prime\prime}=j}=\text{ReLU}[1-(j-j^{\prime\prime})^{2}],$
which allows $f_{i,j}$ to be computed by a constant-width MLP via Lemma C.1.
Finally, for undefined values, $l$ and $f$ are set to $-1$ in the MLP stage using conditional selection (Lemma C.3) based on positional embedding information. In summary, the embeddings generated in this block are ${\bm{e}}^{1}=(l,f)$, and these embeddings are concatenated with the original input embeddings.
Block 2. The second block of the Transformer is designed to compute the first three lines of Algorithm 2. Specifically, this block utilizes the attention mechanism to aggregate the adjacent $m$ bits and derive $s_{i}$. For each token $a_{i,j}$, let $t_{i,j}$ denote the number of preceding tokens (including $a_{i,j}$ itself) $a_{i^{\prime},j^{\prime}}$ such that $\lfloor\frac{j}{m}\rfloor=\lfloor\frac{j^{\prime}}{m}\rfloor$. This block computes the following values:
@@
    $\frac{1}{t_{i,j}}$, where $t_{i,j}$ is as defined above. If $t_{i,j}$ is undefined, its value is set to $-1$.
@@
    $c_{i,j}$, the mean value computed over $a_{i^{\prime},j^{\prime}}p^{j^{\prime}\operatorname{mod}m}$ for previous tokens (including $a_{i,j}$) $a_{i^{\prime},j^{\prime}}$, where $\lfloor\frac{j^{\prime}}{m}\rfloor=\lfloor\frac{j}{m}\rfloor$. If this value is undefined, it is also set to $-1$.
Using these, we derive $s_{w}=c_{i,mw}t_{i,mw}$, where $i$ is the largest index such that the length of ${\bm{a}}_{i}$ exceeds $mk$.
To compute the first value, the query, key, and value vectors are defined as follows:
@@
Using Lemma C.1, the components of the query and key can be computed by preceding MLP layers. The result of the dot product is given by: $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=-\left(\lfloor\frac{j^{\prime}}{m}\rfloor-\lfloor\frac{j}{m}\rfloor\right)^{2}$. This result implies that $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=0$ when $\lfloor\frac{j^{\prime}}{m}\rfloor=\lfloor\frac{j}{m}\rfloor$, and $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle\leq-1$ otherwise. By the definition of $f_{i,j}$ and Lemma C.6, the output of the attention mechanism is $\frac{1}{t_{i,j}}$, as required.
To compute the second value, we redefine the query, key, and value vectors as follows:
@@
Similar to the computation of the first value, we use Lemma C.1 to compute the components of the query and key. In this case, the dot product is given by: $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=-\left(\lfloor\frac{j^{\prime}}{m}\rfloor-\lfloor\frac{j}{m}\rfloor\right)^{2}$. Thus, $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=0$ when $\lfloor\frac{j^{\prime}}{m}\rfloor=\lfloor\frac{j}{m}\rfloor$, and $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle\leq-1$ otherwise. By applying Lemma C.6, the attention output is $c_{i,j}$, as required.
Finally, for undefined values, we assign $-1$ during the MLP stage by employing conditional selection, as outlined in Lemma C.3, utilizing information encoded in the positional embeddings.
In summary, the new embeddings generated in this block can be expressed as ${\bm{e}}^{2}=\left(\frac{1}{t},c\right)$. These embeddings are subsequently concatenated with the original embeddings.
@@
Block 3. The third block of the Transformer computes the value of $c_{i,j}t_{i,j}$. This is achieved by first determining $t_{i,j}$ via the attention layer and $\frac{1}{t_{i,j}}$ from the previous block. Subsequently, $c_{i,j}t_{i,j}$ is computed using Lemma C.1.
@@
Notice that $t_{i,j}$ does not exceed the absolute positional value of the current token. We define the query, key, and value vectors as follows:
  • Query: ${\bm{q}}_{i,j}=\left(\frac{1}{t_{i,j}^{2}},-\frac{2}{t_{i,j}},1\right)$
  • Key: ${\bm{k}}_{i^{\prime},j^{\prime}}=\left(\text{ape}_{i^{\prime},j^{\prime}}^{2},\text{ape}_{i^{\prime},j^{\prime}},1\right)$
  • Value: ${\bm{v}}_{i^{\prime},j^{\prime}}=(\text{ape}_{i^{\prime},j^{\prime}})$
These vectors can be constructed using Lemma C.1. It follows that the inner product
$\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=-\left(\frac{\text{ape}_{i^{\prime},j^{\prime}}}{t_{i,j}}-1\right)^{2},$
@@
which implies $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=0$ if $\text{ape}_{i^{\prime},j^{\prime}}=t_{i,j}$ and $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle\leq-\frac{1}{n^{2}k^{2}}$ otherwise, given that $t_{i,j}\leq nk$. By leveraging Lemma C.6, the attention output is confirmed to be $t_{i,j}$, as required.
@@
Finally, the computation of $c_{i,j}t_{i,j}$ is performed via the subsequent MLP layer. In summary, the embeddings generated in this block are represented as ${\bm{e}}^{3}=(ct)$. These new embeddings are concatenated with the original embeddings to produce the final output of this block.